OpenStudy (anonymous):

What is the value of c such that: x2 + 14x + c, is a perfect-square trinomial?

OpenStudy (anonymous):

Take 1/2 of the coefficient of x, and square the number. The coefficient of the x term is 14. So 1/2 of 14 is 7. And 7^2 = 49. c = 49

OpenStudy (anonymous):

KK Thanks.

OpenStudy (anonymous):

And it checks, because if c = 49 you have x^2 + 14x + 49 = (x+7)(x+7) or (x+7)^2 which is a perfect square trinomial.